### v2

# Context-Aware Code Generation Framework for Code Repositories: Local, Global, and Third-Party Library Awareness

Abstract.
Code generation tools are essential to help developers in the software development process. Existing tools often disconnect with the working context, i.e., the code repository, causing the generated code to be not similar to human developers. In this paper, we propose a novel code generation framework, dubbed $A^{3}$-CodGen, to harness information within the code repository to generate code with fewer logical errors, code redundancy, and library-related compatibility issues. We identify three categories of representative information for the code repository: local-aware information from current code file, global-aware information from other code files, and third-party-library information. Results demonstrate that by adopting the $A^{3}$-CodGen framework, we successfully extract, fuse, and feed code repository information into the LLM, generating more accurate, efficient, and highly reusable code. The effectiveness of our framework is further underscored by generating code with a higher reuse rate, compared to human developers. This research contributes significantly to the field of code generation, providing developers with a more powerful tool to address the evolving demands in software development in practice.
Code Generation, Code Evolution, Code Repository Knowledge Mining
copyright: none
1. Introduction
In the realm of Software Engineering, the advent of Large Language Models (LLMs) has revolutionized various tasks, including code generation [1, 2, 3, 4, 5]. The automatic generation of code by these models significantly reduces development time and effort [6]. Specifically, in the process of evolving software projects, software engineering developers are frequently tasked to add new features and functionality [7]. LLMs such as ChatGPT [8], GitHub Copilot [9], Codex [10] are capable of automatically generating code based on the developers’ requirements (e.g., natural language descriptions of functions and function definitions).
A significant gap still persists between the code generated by LLMs and that crafted by human developers. Typically, when adding new functionality or features to a specific module, developers already had sufficient knowledge about the current task and the entire code repository. Also, according to a previous study [11], developers often expect that code generation tools could be aware of more context/knowledge by using specific frameworks/libraries [12]. In addition, they hope the generated code maximum utilizing specific third-party libraries (e.g., invocation of an API in a library) [13, 14, 15]. Above all, in contrast to the straightforward LLM-based code generation, developers especially take into account three primary categories of information:
1
Local Information: This category encompasses information within the current working context, including defined function signatures and variables. It also covers the relative file paths of the module within the project, such as the module’s Fully Qualified Names (FQNs), like unstructured.documents.html. local information is crucial to ensuring the generated code snippet utilizes accurate variables and operates harmoniously with the other functions in the project.
2
Global Information: This category involves functions from other code files within the same repository, accessible via fully qualified import API names. A previous research highlighted that only around 30% of methods operate independently in open-source projects [16]. Hence, these external functions are essential for implementing new functions without duplicating existing code. For example, developers import functions “from unstructured.partition.text_type import is_bulleted_text” from other modules, instead of writing a function to judge “bulleted text” from scratch.
3
Third-Party-Library Information: This category includes third-party libraries that are already pre-installed in the code repository’s environment. These libraries can be accessed directly through import statements, enhancing the functionality without additional dependencies.
Unfortunately, none of existing code generation methods take above three information into consideration simultaneously. As the most common code generation tool, ChatGPT [8] showcases a commendable level of coding ability [17], but it solely depends on the users’ input. GitHub Copilot [9], which explicitly utilizes local information spanning the current working context, often struggles to harness the full potential of global information presented in the entire code repository. When these methods disregard local and global module information, they risk omitting crucial logic during code generation [18]. For instance, they might forget to extract text from a tag before attempting to identify it as a bullet point (is_bulleted_text). Furthermore, neglecting third-party library information may cause the LLMs to randomly select third-party libraries and generate code based on them. If the selected libraries third-party are unsupported under the given environment, it can lead to compatibility issues, such as “ModuleNotFoundError: No module named xxx.
Therefore, to encounter the insufficient information awareness issue in code generation tasks, we introduce novel $\mathbf{A^{3}}$-CodGen (Local Aware, Global Aware, and Third-Party-Library Aware). This framework systematically mines and utilizes external knowledge from the local module, global module, and third-party libraries. First, our framework begins by acquiring foundational information about functions and third-party libraries within the code repository. This initial step provides a solid function base and a third-party-library base that supports subsequent information retrieval processes. Next, it conducts knowledge retrieval from the function base, the third-party library base, and the current file. These retrieved insights are systematically organized into three distinct categories of code repository-aware knowledge. Subsequently, the framework seamlessly fuses these three types of code repository-aware knowledge into a cohesive prompt. This prompt serves as the input to the foundational LLM, enabling the model to generate code that closely resembles the work of human developers.
Our experiments are driven by two primary objectives: 1) To investigate the effectiveness of three distinct types of code repository information in assisting the model with code generation for specific code repositories. 2) To assess whether our framework can efficiently extract and utilize these three types of information. We crawl and filter 29 representative Python code repositories from PyPI and randomly sample 383 representative functions with comments to create our benchmark dataset. Our experimental results confirm the efficacy of incorporating three types of code repository-awareness knowledge in aiding the foundational model to reuse code repository knowledge. The $A^{3}$-CodeGen framework excels in recognizing and utilizing these three types of code repository knowledge, leading to exceptional performance metrics in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding code reuse capabilities, our framework is capable of generating concise code that seamlessly integrates with the code repository’s working environment. Furthermore, we identify optimal configurations for local and global information, enabling the model to effectively leverage essential knowledge while avoiding the potential pitfalls of information overload.
The main contributions of this paper are as follows.
  • [leftmargin=*]
  • To the best of our knowledge, we are the first to conduct an extensive empirical study, systematically investigating the usefulness of local information, global information, and third-party-library information in code generation tasks.
  • We introduce a novel framework, dubbed $\mathbf{A^{3}}$-CodGen Framework, effectively knowledge mining and knowledge utilization, aiding models in better understanding and fully utilizing the knowledge within code repositories.
  • We constructed a new benchmark dataset, namely RepoEval, to reflect the effectiveness on code-repository-aware code generation. Results demonstrate the validity of the three types of code repository information and show that our framework can effectively mine and utilize this knowledge.
2. Motivation
Figure 1. Knowledge-Starved Code Generation vs. Code Repository Aware Code Generation.
In this section, we discuss the motivation of this study.
2.1. Limitation Of Existing Code Generation Approaches
Benefit by the rapid development of Transformer structure and pre-trained LLMs in the early 2020s, code-LLMs also thrived significantly. Encoder-decoder Transformer code models were dominant, such as the milestone work CodeT5 [19]. Also, other code models like CodeBERT [20] adapt the original BERT architecture to code generation tasks, seamlessly bridging natural language understanding with code comprehension. Building on this, GraphCodeBERT [21] extend its capability by incorporating the graph structure of code to enhance performance. Whilst, more recently, the decoder-only Transformer models have become the mainstream of generative models. Codex [22] and CodeGen [23] are representatives, and many others, like [24, 25, 26, 27, 28], also belong to this category.
Nowadays, to Software Engineering developers, the LLM-empowered code generation has become an irreplaceable assistant tool in their development cycle [29, 30]. Although the code-LLM has been developed for years, those code generation tools still play a supporting role in the actual SE development. Essentially, there are significant differences between the code generated by LLMs and that crafted by human developers. Specifically, we conclude this phenomenon into the following three aspects:
  • Observation 1: Followed by the innate features of LLMs, the content and quality of generated codes heavily depend on the human-crafted prompts.
    Discussion 1: Providing longer prompts with richer content seems a plausible solution. With the growth of PLM size, the maximum input length increases from 512 tokens for the early models [31] to 32,000 tokens for the extended GPT-4 models [32]. However, it is still impossible to feed the entire code repository and all other relevant information into the LLMs. Furthermore, studies have demonstrated that irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performance [33]. Therefore, to maintain the practicability and effectiveness simultaneously for code generation tasks, external context information should be extracted and selectively fed into LLMs.
  • Observation 2: Existing methods focus on generating independent and often small-scale code units [34, 16]. According to a previous study [34], the code generated by LLMs shows lower method-level correctness. Although various benchmarks have been proposed (Details are discussed in the Related Work (Section 7)), newly generated code still remains questionable.
    Discussion 2: The correctness relates to many factors, and evaluation of correctness is already challenging enough. Studies have shown that larger and more powerful foundation LLMs lead to better code generation correctness. Additionally, retrieving and reusing existing functions or APIs (which are usually verified by other human developers), as the skeleton for code generation, could also fundamentally improve the correctness. This is also the strategy that human developers commonly apply in practice.
  • Observation 3: Some code-LLMs (e.g., GPT-4) have developed superior skills in simple algorithms such as greedy algorithm and searching strategies, but they perform poorly in complicated data structures such as tree structure and graph theory [35].
    Discussion 3: Similarly, instead of generating complex code snippets from scratch, importing functions that are already implemented can mitigate the shortcomings of advanced-level programming challenges. For example, Python developers are likely to import functions from networkx [^1] https://networkx.org/documentation/stable/index.html package to build a graph, not building graphs by the dictionary-like data structure.
2.2. Code Generation in Code Repository
During software development, programmers typically work within a specific code environment, which is essentially a designated code repository. This code repository encompasses all the project’s modules, functions already present in these modules, and third-party libraries. It serves as the foundational framework upon which developers build and expand their software applications. In this context, developers (especially experienced ones) must navigate and understand this code repository environment thoroughly to ensure seamless integration of newly added code with the existing one. When it comes to LLM-based code generation techniques, solely relying on inputting developer requirements to these models fails to fully utilize the potential of LLMs in interacting with existing code repository environments. Specifically, those techniques are blind to the repository’s unique naming conventions and programming styles when the completion target is relevant to cross-file context [36, 37, 38]. Therefore, instead of using a single code file, well-designed programs emphasize separating complicated functionality into independent modules, each implementing only one aspect of the desired functionality [39, 40].
[…]
Furthermore, since the model is unaware of the existence of the ”check text for bullet characters” function in the global module, it defines a new function is_bullet_text (highlighted in the blue box) in the generated code. A large part of this code duplicates the function is_bullet_text in the global module unstructured/partition/text_type.py, resulting in code redundancy. It also brings additional concerns about the correctness of the newly generated function is_bullet_text.
Additionally, since the model is unaware of which third-party libraries are available, it arbitrarily imports and uses the bs4 (BeautifulSoup4222https://pypi.org/project/beautifulsoup4/) library. This can result in the generated code that refers to a library that is not available, leading to the bug “ModuleNotFoundError: No module named ***”.
These knowledge gaps highlight the limitations of existing code generation approaches, emphasizing the need for more context-aware methods that can effectively interact with the code repository environment. In summary, a lack of knowledge about local and global modules can result in models that generate code that lacks functional constraints or contains redundancies; and a lack of knowledge about third-party libraries in the code repository can lead to models using arbitrary third-party libraries, which can cause problems with the generated code not working. As a result, generating code in specific modules of the code repository is more than just meeting functional requirements. To generate code similar to human developers, the approach must fully understand and effectively utilize this diverse contextual knowledge to ensure that newly introduced code is seamlessly coordinated with the existing code repository.
2.3. Bridging the Gap: Making LLMs Aware of Three Types of Repository Knowledge
Above all, in order to enable LLMs to fully utilize information spanning over the code repository environment for code generation, it is essential to provide the model with three types of knowledge: local modules, global modules, and third-party library information.
As shown in Fig. 1 Local Module Aware & Global Module Aware & Third-Party Library Aware Code Generation, when the developer requirement is entered, since the model is aware of the knowledge of the local module, the generated code reuses the _construct_text function (highlighted in yellow) in the local module to extract the text content in the text label element. In addition, because the model is aware of knowledge of the global module, the generated code reuses the is_bulleted_text function (highlighted in blue) from the global module unstructured/partition/text_type.py, thus reducing redundant code definitions. In addition, the model is aware of the existence of third-party libraries and therefore reuses the lxml library to ensure its availability.
In the next section, we provide a comprehensive description of how to extract, structure, and utilize local aware, global aware, and third-party library aware information from code repositories.
Figure 2. $A^{3}$-CodGen Framework.
3. $A^{3}$odGen Framework
In this section, we specifically elaborate on our novel $A^{3}$-CodGen Framework to harness the full potential of code repository knowledge for assisting LLMs in code generation.
As depicted in Fig. 2, the $A^{3}$-CodGen Framework comprises three essential phases. First, we obtain the function base and third-party library base in the code repository knowledge construction phase (Section 3.1), and they provide sufficient knowledge to support subsequent knowledge retrieval. Then, we retrieve the code repository knowledge from other code files, the third-party library repository, and the current file. After that, we organize them into 3 types of code repository aware knowledge (Section 3.2). Finally, in the knowledge utilization phase (Section 3.3), the three code repository-aware knowledge are fused into a prompt, which serves as an input to the foundation LLM, to ultimately generate codes similar to human developers with three essential awareness.
3.1. Repository Knowledge Base Construction
Code repositories contain a wealth of knowledge to exploit, but this knowledge usually exists in an unstructured form and cannot be directly utilized effectively, scattering in the code repository. Therefore, we need to perform systematic knowledge extraction to transform and compress this knowledge into a structured form in order to build a knowledge base to support subsequent knowledge utilization, such as knowledge retrieval. Thus, we perform code repository information extraction (Section 3.1.1) and third-party library extraction (Section 3.1.2) to obtain the function base and third-party-library base, respectively.
Figure 3. A Piece of Code in the Module & Function Summarizer Unit
3.1.1. Code Repository Information Extraction
Given a code repository, we can effortlessly obtain all source code files in the repository’s root folder. Then, for each source code file, we can extract all functions included. For each function, we perform both direct code information collection and semantic code knowledge extraction. In total, by snowballing this process, all essential repository information is extracted.
Basic Code Knowledge Extraction. Abstract Syntax Tree (AST) is a tree representation of the source code, and studies have demonstrated its usefulness on various SE tasks [41]. In this study, we employ the Python library ast [^3] https://docs.python.org/3/library/ast.html to analyze the AST for each source code file in the code repository. From the AST, we extract six essential types of basic information: 1) Comment. For example, in Fig 3, the comment for the function element (circled in the yellow box) is Get all elements from pages in sequential order. 2) Signature. This knowledge illustrates the parameters and returns of a function. For example, in Fig 3, the signature of the function element (circled in the yellow box) is def elements(self) -¿ List[Element]. 3) Function Source code. 4) Class associated with the function. For example, the function element in Fig. 3 corresponds to the class Document. 5) Is_Empty tag. True means empty function, False means empty label. For example, function element has is_Empty value False. 6) The file path of the function. This reflects the relative location of the current function in the entire code repository. For example, the file path of the function element is unstructured/documents/base.py.
Semantic Code Knowledge Extraction. AST is criticized for its prone to the long-term dependency problem. Also, simply employing AST potential weakens the capability to capture real and complex semantics [41]. Thus, we further maneuver Semantic Code Knowledge Extraction to enhance the basic code knowledge with additional semantic information. During this process, we extract three types of semantic code knowledge, as follows:
1) Fully Qualified Name (FQN) of Functions. FQN refers to the complete name of a function, typically composed of the module path, class name (if the function belongs to a class), and the function name. For example, in Fig. 3, the FQN of the function element (highlighted within the yellow box) is unstructured.documents.base.Document.elements. FQN is often used to represent structural information about a function, indicating which module it originates from in the code repository and, if applicable, which class within the module. Developers can import global functions from the global module using their FQNs.
2) Function Summary. Not every function is elaborated by comments. For instance, in Fig. 3 (a), the function __init__ (encircled in a red box) lacks a comment. Even though the comments may exist, their quality varies distinctly. Some do not clearly state the basic functionality and usage of the function; whilst other comments may cover too many non-essential details such as notifications, technical details, or other secondary information unrelated to the overall functionality. Over-length textual descriptions may distract people from understanding the actual usage and operational behavior of the function. Therefore, we generate the function summary to precisely describe the purpose and functionality. Much previous work has shown that LLMs are capable of in-context learning [42, 43, 44], just input a task description and some examples can make LLMs behave well. Thus we use GPT-3.5-Turbo as the foundation model to implement the function summarization task.
Function Summarizer Unit. To generate natural language semantic information for a function, we design a prompt as the core of the Function Summarizer unit. This prompt creates summaries for each function. As depicted in Fig. 3, our Function Summarizer prompt consists of a task description (highlighted in purple), five demonstration examples, and the input template. The input to this unit is the function’s code (highlighted in green), and the output is the model-generated summarization for that function (highlighted in red). For instance, when provided with the code of the __init__ function within the red box in Fig. 3, the resulting summary is Initialize the object with two optional attributes: _pages and _elements.
3) Vector Representations. In the previous two steps, we underlined explictly select information. In order to facilitate subsequent retrieval of global functions, it is crucial to possess an effective code representation that accurately captures the semantics of functions. Therefore, we employ a dual-pronged approach, conducting separate embeddings for both the source code and the function summaries. This results in the creation of code vector and summary vector, respectively.
3.1.2. Third-Party Library Extraction
As third-party libraries commonly appear in code files, it is often the case that when we provide specific third-party libraries to a model, the model is likely to have already learned how to use those libraries during the training process [45, 46]. Therefore, we only need to provide the name of the library and the model will be able to automatically understand how to use that library to perform a specific task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The generated code example is as follows: $cosine\_similarity=numpy.dot(vector\_a,vector\_b)/(np.linalg.norm(vector\_a)*numpy.linalg.norm(vector\_b)).$ By simply providing the name of the library and the task requirements, the model can generate the appropriate code, helping the developer accomplish programming tasks quickly. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extracted the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base.
3.2. Three Types of Repo-Aware Knowledge Retrieval
As we discussed in Section 2.1, although the maximum input length grows greatly, it is still impossible to feed the entire code repo and third-party library information to LLMs. Furthermore, irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performance [33].
To enable the model to fully utilize the three types of information extracted from the code repository during code generation, we conduct comprehensive knowledge mining: extracting four types of local information (Section 3.2.1), retrieving global functions from the function base using developer requirements to obtain global-aware information (Section 3.2.2), and structuring third-party library information from the third-party library base (Section 3.2.3).
Figure 4. Local-Module-Aware, Global-Module-Aware, Third-Party-Library-Aware Knowledge
3.2.1. Local-Module Knowledge Mining
Within the user’s working context, referred to as the local module, four types of local-module information can be encountered:
  • [leftmargin=*]
  • Local functions. Local functions encompass all functions that are defined and implemented within the local module. We extract these local functions and represent them in the form of ¡Fully Qualified Name (FQN), Function Summary, Function Signature¿. The FQN provides details about the function’s location within the module, the Function Summary reflects the natural language description of the function’s purpose, and the Function Signature reveals its input and output specifications. Fig. 4 (a) illustrates local functions, displaying the FQN, summary, and signature for each function, such as function k. We use indentation to differentiate whether a function belongs to a class. For example, function 1 belongs to the HTMLDocument class, indicated by an indented tab, while function k is a module function not associated with any class. Additionally, each function’s information is enclosed within {} braces. This structuring serves the purpose of organizing the information for all local functions, aiding the model in more effectively recognizing local function knowledge.
  • Class Instance Attributes. Class instance attributes, also known as instance variables, are variables defined within a class’s constructor function (__init__ in Python). These variables are specific to instances of the class and store instance-specific attributes or state information. Some local functions may be defined within a class, and they might involve interactions with class instance attributes, accessing or modifying them to achieve specific functionality. For instance, in Fig. 4 (a), the source code within function1 explains how class instance attributes are defined. Therefore, to represent class attributes, we do not extract them separately but include the source code of the constructor function (__init__) that defines these class attributes directly.
  • Local Module FQN. Local Module FQN partially reflects the structural information of the working context (local module) within the entire code repository. It helps the model understand the context of how local functions, class instance attributes, and variables within this module are used. We derive the Local Module FQN from the current file path, providing a contextual reference for the code within this module. For example, the local module FQN of the file in Fig. 4 (b) is unstructured.documents.html.
  • Module variables. Local modules may define variables, some of which can directly interact with local functions, potentially supporting the functionality of newly added functions. We unify the declarations of these variables. However, if there are no variables defined within the module, we indicate this with None.
After obtaining these four types of local-module knowledge, we organize this knowledge in a structured format. For instace, as illustrated in Fig. 4 Local Module Aware Knowledge.
3.2.2. Global-Module Knowledge Mining
In addition to using local module knowledge, global module knowledge also plays a significant role in code generation. Instead of letting the model implement complicated logic steps from scratch, we expect the model can directly call functions if they are already present in other code files. To help models realize that global functions can be reused, we retrieve global functions in the code repository and organize them in a structured way. The global module knowledge mining process is divided into three phases: what-if code generation, global function retrieval, and information unification.
Figure 5. Two Types of Code Generator Units
  • [leftmargin=*]
  • What-If Code Generator Unit : As illustrated in Fig. 5 (a), in the task description, we use a structured prompt that defines roles, task requirements, and output requirements for the LLM. In the prompt, we define two terminologies: developer_requirement, and to_be_generated_code. The developer_requirement contains the function description and function definition. In task description part (highlighted in purple), we set three rules and use commands to guide the model make chain-of-thought. Then, we provide three representative examples for in-context learning. The inputs is the developer_requirement (highlighted in green), and output is the to_be_generated_code (highlighted in red). Notably, the generated code is only used to retrieve the global function, so we do not require its completeness as it is not the final code. This helps the retrieval of relevant reusable functions in the next Global Function Retrieval Unit.
  • Global Function Retrieval Unit: This unit uses function descriptions and the what-if function-code as input to retrieve reusable global functions from the function base. Specifically, we first embed function_description into $vector_{des}$ and compute the cosine similarity between this vector and the summary vectors of all functions in the function base. Thus, we can rank and obtain the top k $RetrievedFun_{des}$. Second, we embed the what-if function-code as a $vector_{f}$ and compute the cosine similarity between this vector and the code vectors of all the functions in the function base, so we can rank and obtain the top k $RetrievedFun_{code}$. Finally, we merge the sets of $RetrievedFun_{des}$ and $RetrievedFun_{d}es$ to obtain the global function that is most relevant to the provided input, which is denoted as $RetrievedFun$
  • Information Unification Unit: Above all, for each $RetrievedFun$, we organize and display the retrieved information as ¡FQN, function signature, summary¿ in a structured format. Fig. 4 (c) presents the global-aware knowledge after unification. Similar to local functions, the information for each function is enclosed within {} braces to assist the model in more efficiently recognizing function knowledge.
3.2.3. Third-Party Library Knowledge Mining
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in section 3.1.2. It’s important to note that we extract only the names of these libraries and do not include any of their code. Third-party libraries are publicly accessible and widely applied, so according to our empirical observation, we found that the model has already encountered and learned most of the libraries during its training process. Thus, we only extract the names of these libraries without any of their code. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
3.3. $A^{3}$ode Generation
After section 3.2, we obtain structured local-aware knowledge, global-aware knowledge, and third-party library-aware knowledge. To ensure the model’s full awareness of these three types of knowledge, we design an $\mathbf{A^{3}}$ Code Generation Unit that seamlessly integrates user requirements with the three knowledge sources to generate $A^{3}$ functional code.
The $A^{3}$ Code Generation Unit, as depicted in Fig. 5 (b), consists of four key components: task description, examples, input, and output. In the task description, we guide the model through the process of code generation step by step, employing a chain-of-thought approach [47, 48, 49].
  • [leftmargin=*]
  • Firstly, in @command1, we prompt the model to contemplate the logic and steps required to fulfill the developer’s request.
  • Next, in @command2, we encourage the model to consider whether there is any reusable information within the local-aware knowledge and utilize it if available.
  • Following that, in @command3, we guide the model to think about whether there are any reusable global functions within the global-aware knowledge and utilize them if present.
  • Subsequently, in @command4, we prompt the model to explore how to leverage the third-party library-aware knowledge to interact with libraries effectively.
  • Finally, in @command5, if the model cannot leverage any information, it is instructed to generate code directly based on the logic and steps it has formulated.
  • Additionally, we have included an @command that reiterates the importance of strictly following the steps from 1 to 5.
We establish three rules to govern the code generated by the model, which include calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as occasionally, the model may generate abstract functions) We then provide 3 representative examples. Finally, we input our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig. 5 (b)).
4. Experiment Setting
In this section, we describe our dataset and evaluation metrics used to estimate the performance of our framework. We select GPT- 3.5-Turbo as the default foundation LLM for this study.
We propose the following four Research Questions (RQs):
  • [leftmargin=*]
  • RQ1: Does Local Module Aware Knowledge indeed help LLMs fully utilize local module knowledge during code generation, and which types are most effective in improving local awareness?
  • RQ2: Is Global Module Aware Knowledge effective in assisting LLMs to utilize global module knowledge during code generation, and what is the optimal “k” value for retrieving global functions?
  • RQ3: Does Third-Party Library Aware Knowledge significantly impact LLMs’ awareness of available third-party libraries during code generation?
  • RQ4: What is the performance of our $A^{3}$-CodGen framework?
4.1. Dataset Collection
While there exist some benchmark datasets designed for code generation, such as [50, 51, 34, 52, 16], They suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilities. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. Therefore, to address the potential risk of data leakage and accurately reflect the code generation capability, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify whether the models truly grasp $A^{3}$ code generation after receiving the information we provide.
4.1.1. Code Repository Collection
According to OpenAI’s official website [^4] https://platform.openai.com/docs/models/gpt-3-5, the training data for GPT-3.5-Turbo is cut off as of September 2021. To ensure that our benchmark dataset contains code repositories that have not been seen during model training, we select projects with the earliest release date after October 1, 2021 on PyPI. We further filter the code repositories by the following criteria: The corresponding GitHub code repositories have more than 100 stars and 100 forks. As the number of stars and branches approximately indicates the quality of the repository. We further selected repositories with more than 10 code files to ensure they have a sufficient number of global modules to help create the benchmark dataset. Ultimately, we collected 29 code repositories that meet these criteria for creating the benchmark.
Then, we constructed 29 function bases and 29 third-party library bases for these 29 code repositories following the method described in Section 3.1. In the function bases extracted for these 29 code repositories, there are a total of 13,784 functions. Among these functions, 10,077 are class member functions, while the remaining 3,707 are standalone functions defined within modules. Since class member functions often involve complex relationships related to inheritance and other classes, it becomes challenging to analyze the code invocation relationships. Therefore, our benchmark dataset in this study focuses solely on standalone functions. From these 3,707 standalone functions, we filter out functions that have comments and are non-empty. In total, we obtained 1,533 functions that meet these criteria.
4.1.2. Benchmark Construction
Given the comments and the definition of a function, the corresponding source code for that function serves as the standard answer to be generated. Simultaneously, we extract call information for the function to determine the extent of code reuse within it, because a higher extent of function reuse reflects the framework learns better from three awareness.
Given a function, code reuse falls into one of the three categories: 1) It could be part of locally defined functions within the local module; 2) belong to a global module; or 3) belong to a third-party library. To obtain this reuse information, we need to analyze the call relationships between functions among codes. However, existing code analysis tools like pycallgraph [53] require code compilability to accurately generate call graphs for functions. Considering that it is impractical to set up running environments for all these code repositories due to diverse requirements and potential conflicts, and given that static analysis tools like pyan [54] have relatively lower accuracy and cannot meet our needs, we resort to manual annotation to extract the call information between functions.
The manual annotation for each function is resource-intensive and time-consuming. To ensure the efficiency of our experiments, we employed a statistical sampling method [55]. Specifically, we performed sampling on the 1,533 standalone functions with comments and non-empty code. We ensured that the estimated accuracy within the sample remains within a 0.05 error range at a 95% confidence level while ensuring that functions from each code repository are included in the sample. Finally, we acquired 383 functions, which are used for subsequent experiments.
4.2. Metrics
To comprehensively evaluate whether the three awareness are effective and whether the generated code has successfully perceived three awareness, the following metrics are employed to assess the generated code in different aspects:
  • [leftmargin=*]
  • Precision: This indicates how much of the code generated by the model is reused correctly.
  • Recall: This indicates how much of the code that really needed to be reused was correctly reused by the model.
  • F1 Score:The harmonic mean of Precision and Recall, which measures the overall effectiveness of the model in reusing code.
  • Library Coverage: To evaluate whether the third-party libraries used in the generated code are exist in the code repository’s existing environment, we calculate library coverage. This metric represents the proportion of libraries used in the generated code that are already pre-installed in the code repository’s environment.
  • Lines of Code (LOC): To assess the conciseness of the generated code, we calculate the number of lines of code (LOC). LOC is determined by counting the lines in the code after removing comments. Fewer lines indicate that the generated code is more concise.
  • CodeBLEU [56]: This is an improved version of the BLEU metric, designed for code, and is a weighted average of the lexical, abstract syntax tree, and data flow matches.
Subsequently, we enlist two developers with more than five years of experience in Python programming to conduct independent evaluations. For each generated result, they assess whether the function utilizes local functions, assigning a score of 1 if it does and 0 if it doesn’t. The same process applies to global functions and third-party libraries. Then we calculate the TP, FP, TN, FN, Precision, Recall, F1, and Accuracy for the local functions awareness. After completing their evaluations, they discuss any discrepancies and reach a consensus. After the evaluations, discrepancies are discussed, and a consensus is reached.
In all the upcoming experiments, we use this benchmark for evaluation and utilize the same method to annotate the code generated by the model. The calculated Cohen’s Kappa coefficient [57] for all annotated data across the research questions fall within the range of 86% to 95%, indicating a high level of agreement among the various assessments.
4.3. Foundation Models
As for the LLM selection, we used the gpt-3.5-turbo model, which demonstrates excellent performance in various software engineering tasks. As for the embedding model, we selected the ada-embedding-002 model [58], which has the capability of generating high-quality sentence embeddings, thus enabling us to perform effective similarity computation and retrieval operations. Similar to other LLM-empowered frameworks in SE domain, the more powerful foundation model is selected, the better performance can be achieved, generally.
5. Evaluation
5.1. The Effectiveness of Local-Aware Knowledge (RQ1)
Local Information can assist LLMs in code generation by making them aware of local contextual information. We extract and organize four types of Local-Aware Knowledge from the local module: local functions, class instance attributes, local module FQNs, and module variables. These four types of essential knowledge provide different perspectives on the context of the local module. We aim to investigate whether Local-Aware Information can genuinely help LLMs reuse local functions during code generation. Additionally, we also aim to identify the optimal local knowledge configuration, determining whether all four categories of knowledge should be made available to the model’s awareness or only specific ones.
We use vanilla ChatGPT as a baseline, which does not have any local information in its input. We define four variants of local-aware methods, as shown in Table 1.
  • [leftmargin=*]
  • +Local Functions signifies the use of Local Functions knowledge from the current file context.
  • +Class Instance Attributes indicates the use of Class Instance Attributes Knoweldge
  • +Module Variables involves using the module Variables knowledge in the current file context.
  • +Local Module FQN employs the fully qualified name corresponding to the current file.
We run these five different configurations on the benchmark to obtain code generation results.
Table 1. Comparison of the results obtained with different Local Information. And find the best configuration for Local Information Extraction.
Variant Precision Recall F1 Accuracy
ChatGPT 0 0 0 0.754
ChatGPT + Local Functions 0.5 0.559 0.528 0.780
ChatGPT + Local Functions              +Class Instance Attributes 0.563 0.623 0.592 0.809
ChatGPT + Local Functions              +Class Instance Attributes              +Module Variables 0.489 0.567 0.525 0.783
ChatGPT +Local Functions              +Class Instance Attributes              +Module Variables              +Local Module FQN 0.542 0.629 0.582 0.809
5.1.1. Result Analysis
Table 1 displays the results corresponding to code generation using vanilla ChatGPT and four types of Local-Aware Knowledge Configurations. vanilla ChatGPT has a precision and recall of 0 because it does not recognize any local functions. Consequently, in its generated code, new functions are often defined to complete certain logic, resulting in code redundancy.
When we provide the model with knowledge of Local Functions, it immediately gains the ability to utilize local functions, resulting in a Precision of 0.558, a Recall of 0.624, an F1 of 0.525, and an accuracy of 0.779. This is because knowledge of Local Functions enables the model to directly invoke local functions, reducing code redundancy and enhancing code reusability and efficiency.
Upon additionally providing knowledge of Class Instance Attributes, the model’s Precision, Recall, and F1 score improve by 12.78%, 11.44%, and 12.15%, respectively, compared to providing only knowledge of Local Functions. This improvement is attributed to the provision of Class Instance Attributes, which enables the model to better understand how local functions interact with class variables in the project, leading to more accurate application of local functions.
However, when we introduce the Module Variables, the model’s performance drastically declines, particularly with a 13.21% drop in Precision. This may be due to the increased information load resulting from knowledge of Module Variables, which negatively impacts performance.
Nevertheless, the model’s performance is restored and approaches the capabilities of providing knowledge of Local Functions and Class Instance Attributes when we subsequently add knowledge of Local Module FQN. This suggests the existence of subtle interactions within the knowledge configuration. When the model receives only one or two key knowledge types, its performance may be superior to when it receives more information. However, as the volume of knowledge increases to a certain extent, the model’s capabilities strengthen, resulting in improved performance.
Ultimately, we opt for the Local-Aware Knowledge configuration, which includes knowledge of Local Functions and Class Instance Attributes. This configuration will be used as the default in the following experiments.
Finding 1: The Local-Aware Knowledge significantly enhances the model’s ability to reuse local functions, thereby reducing code redundancy. The best configuration is the combination of Local Functions and Class Instance Attributes.
5.2. The effectiveness of Global-Module-Aware Knowledge (RQ2)
In this RQ, we aim to investigate whether global-aware information can genuinely help LLMs reuse local functions during code generation. Additionally, we seek to determine the optimal top-k value for retrieving candidate functions, obtaining the best global-aware configuration.
Specifically, we employ No Global-Aware as the baseline and this involves providing only Local-Aware Knowledge without Global-Aware Knowledge input. Furthermore, to determine the optimal top-k value, we define four variants, namely, k=1, 5, 10, and 15. We run the total ive different configurations on the benchmark to obtain code generation results. Subsequently, we assess the generated code using the same method as in section 5.1, where each result is evaluated for whether it uses global functions.
Table 2. Comparison of the results obtained for different Top k numbers of retrieved functions to find the best k value. Where Precision, Recall, and F1 are for global functions.
Global functions of different top k Avg. Retrieved Functions Precision Recall F1 Accuracy
No Global-Aware (k = 0) 0 0 0 0 0.819
k=1 1.757 0.535 0.441 0.484 0.788
k=5 8.154 0.658 0.533 0.589 0.825
k=10 16.031 0.550 0.441 0.490 0.793
k=15 23.592 0.521 0.418 0.464 0.783
5.2.1. Result Analysis
Table 2 presents the results obtained using ”No Global-Aware” and four different top-k values for retrieving global functions. The No Global-Aware has both precision and recall as 0 because of lacking of global-aware information as we expected. Consequently, the generated code only utilizes local functions to accomplish tasks, and it resorts to defining new functions for specific logic if local functions are insufficient. When k=1, the model shows some awareness of global functions to a certain extent, underlined by the precision and recall as 0.535 and 0.442, respectively. At k=5, the model’s performance significantly improves, indicating that the model now has a more comprehensive understanding of global-aware information. However, when k=10, the model’s F1 score drops by 16.76%, indicating a substantial reduction in the model’s ability to reuse global functions. As k rises to 15, the effectiveness of the model decreases further. This is because the overdosed amount (between 10 to 20 functions) of functions is retrieved, introducing excessive interference and hindering accurate reusing.
To further assess the impact of the number of global functions presented to the model, we compute the average number of retrieved functions, denoted as Avg. Retrieved Functions. We find that providing the model with around 8 functions yields the best performance in terms of code reuse, striking a subtle balance. Thus, we select k=5 as optimal configuration for subsequent experiments.
Finding 2: Global-Aware Information significantly enhances the model’s ability to reuse functions in other code files in the repository. The optimal configuration for retrieving global functions is k=5.
5.3. The effectiveness of Third-Party-Library-Aware Knowledge (RQ3)
In section 3.2.3, we delve into Third-Party-Library Aware information, which helps the model recognize pre-installed third-party libraries in the code repository. To validate that point, we compare code generation between with and without Third-Party-Library Aware information. Only Local-Aware and Global-Aware information is fed into the model. Then, we run these two different configurations on the benchmark to obtain code generation results. Notably, we introduce the Library Coverage, the percentage of third-party libraries used in the generated code in the third-part-library base, to further elaborate the extent of reusing.
Table 3. Compare the difference between having Third-Party-Library Aware and not having it on Precision, Recall, F1, Library Coverage.
Third-Party-Library Awareness Precision Recall F1 Accuracy Library Coverage
No 0.464 0.620 0.530 0.676 0.686
Yes 0.599 0.901 0.719 0.815 0.906
5.3.1. Result Analysis
Table 3 demonstrates that after incorporating Third-Party-Library Aware Knowledge, the model exhibits significant improvements in precision, recall, F1 score, and accuracy, with increases of 29.14%, 45.44%, 35.66%, and 20.47%, respectively. This enhancement can be attributed to the fact that, with Third-Party-Library Aware information, the model is suggested to use a specific subset of libraries, narrowing down its choice range and making it more likely to generate desired code. Our finding is further proven by that methods with Third-Party-Library Aware information outperform those without such knowledge in the Library Coverage metric by 32.09%. However, library coverage does not reach 100% due to occasional instances where the model obstinately believes that there are no viable libraries in the third-party repository to implement the code, which could be attributed to model hallucinations [59, 60, 61].
Finding 3: Third-Party-Library Aware information significantly improves the model’s ability to utilize available third-party libraries, reducing the risk of potential compatibility issue.
5.4. The Effectiveness of $A^{3}$odGen Framework (RQ4)
In previous sections, we have individually verified the effectiveness of Local-Aware, Global-Aware, and Third-Party-Library-Aware information. We also explored the optimal configurations. In this section, we aim to integrally validate the effectiveness of our $A^{3}$-CodGen Framework from various perspectives. We compare our $A^{3}$-CodGen with three baselines: vanilla ChatGPT, Local-Aware Generation, and $A^{2}$ Generation (Local-Aware and Global-Aware). In addition to metrics based on the confusion matrix, we also calculate library coverage, line of code (LOC), and CodeBLEU. Among these metrics, a lower LOC indicates simpler code, and a higher CodeBLEU suggests that the generated code closely resembles the ground truth.
Table 4. Comparison of the overall effect of Framework with Ablation Baselines
Metrics
Vanilla
ChatGPT
Local-Aware
Local-Aware
+Global-Aware
Local-Aware
+Global-Aware
+3rd-Party-Lib-Aware
Local Reuse Precision 0 0.564 0.719 0.701
Recall 0 0.624 0.570 0.654
F1 0 0.592 0.636 0.677
Accuracy 0.755 0.810 0.794 0.830
Global Reuse Precision 0 0 0.658 0.648
Recall 0 0 0.533 0.605
F1 0 0 0.589 0.626
Accuracy 0.820 0.820 0.825 0.856
Third-Party Library Reuse Precision 0.8 0.687 0.464 0.599
Recall 0.492 0.563 0.620 0.901
F1 0.609 0.619 0.530 0.719
Accuracy 0.598 0.668 0.676 0.815
Library Coverage 0.792 0.855 0.686 0.905
Avg Line of Code (LOC) 16.320 16.487 12.598 12.804
CodeBLEU 0.339 0.344 0.312 0.322
5.4.1. Result Analysis
As shown in Table 4, our $A^{3}$-CodGen performs exceptionally well in terms of Local Reuse, Global Reuse, and Third-Party-Library Reuse, demonstrating its capability to effectively leverage information from the code repository. Furthermore, the framework generates code with optimal line of codes (LOC), indicating its capability to produce concise and efficient code.
Regarding CodeBLEU, we observed an interesting phenomenon: as our model’s code reuse capability increases, the value decreases. After manual inspection, we found that the unwanted excessive code reuse led to the decline of CodeBLEU. Specifically, as our model tends to make full use of existing code and functions by design, it generates distinctly different but more concise code compared to ground truth. Based on our manual evaluation, we surprisingly found that a large amount of generated code is better than the ground truth, i.e., our generated code subtly reuses existing functions, making the code more efficient, readable, and maintainable. For example, the function named allcap_differential_getter determines whether the input span contains mixed-case words. In the ground truth, a longer code snippet is written to accomplish this task (10 lines), but our approach generates a shorter, more streamlined code snippet that achieves the same goal by invoking a global function src.asent.getters.__get_lower_txt in another code file. This further underscores the effectiveness of our framework, which not only generates code similar to human developers but also assists developers in recommending more reusable code snippets.
Finding 4: $A^{3}$-CodeGen framework excels in awaring three types of code repository knowledge, resulting in an excellent performance in terms of local reuse, global reuse, and third-party library reuse. Thanks to its outstanding reuse capabilities, our framework can generate concise code that aligns with the code repository’s working environment. $A^{3}$-CodeGen maximizes code reuse, and to some extent, the quality of the generated code surpasses that of the ground truth.
6. Discussion
New Paradigm for Code Generation Tasks. Previously, code generation tasks aimed to automate code production to meet specific requirements, and the advent of LLMs like GitHub CoPilot and ChatGPT has substantially mastered this task. However, practical code generation tasks in real-world scenarios extend beyond mere requirement satisfaction. Generated code is expected to be aligned with project-specific requirements and leverage existing knowledge simultaneously, leading to better readability, maintainability, and practicability, which can further help developers in the software development process. Therefore, we hope this work sheds a light on this emerging but significant code generation topic: generating context-aware code that is compatible with the working environment.
Knowledge Engineering for Code Generation. LLMs are sensitive to prompts and well-designed prompts are decisive in obtaining high-quality results [62, 63, 64]. Either unorganized or too lengthy prompts are impractical. Instead, it is advisable to extract semantic-enriched knowledge from the code repository to guide the model to directional generate code. However, how to selectively fuse the different types of knowledge presents a challenge, potentially impacting the model’s awareness of different knowledge types. In Section 5.1.1, we find that providing diverse information does not necessarily improve model performance. Meanwhile, in section 5.4, we find that providing the model with global awareness and third-party library information unexpectedly enhances its ability to reuse native functions. This highlights the interplay between different types of code knowledge and emphasizes the importance of effectively mining and leveraging code knowledge from libraries in future research.
Threats to Validity. The threats to validity of this paper lies in three aspects: 1) Subjective Labeling. The data used in our study involves manual labeling, which introduces subjectivity in the annotations. We take a rigorous approach by enlisting the expertise of two professional developers to label the code. Furthermore, any instances with ambiguous labels are discussed and resolved to establish a reliable ground truth.
2) Limited Dataset Selection. We conduct experiments using a selection of 29 Python code repositories. While this number may appear limited, it’s essential to note that the repositories are carefully chosen based on their high quality, as indicated by a substantial number of stars and forks. These repositories are considered representative of the Python programming ecosystem. Additionally, to ensure the credibility of our ground truth, we employ rigorous sampling techniques to maintain a confidence level of 95% or higher. 3) General Applicability of Our Framework. Our framework is generic in the sense that its design is based on generalized code generation principles and natural language processing techniques, and does not depend on the specific nature of a particular programming language or domain. Although we chose Python as an example in our study, this was done only to validate the effectiveness of the framework.
7. Related Work
Code generation tasks is about creating source code from provided natural language descriptions or requirements, which has been a longstanding focal point in software engineering research.
Repository-Level Code Completion. Compared to traditional function-level code completion, repository-level code completion has drawn attention to practical developments [65, 38, 36, 66]. Shrivastava et al. generated example-specific prompts using a prompt proposal classifier [65] and further proposed RepoFusion to incorporate relevant repository context by training code models [38]. In addition, Ding et al. [36] proposed a cross-file context finder tool, named CCFinder, which constructs a project context graph and retrieves the most relevant cross-file context based on local import statements. RepoCoder [66] retrieved similar code snippets from the repository as part of a prompt and proposes an iterative retrieval-generation paradigm. This novel strategy bridged the gap between the retrieved context and the intended target through an interactive process.
Retrieval Augmented Code Generation. Retrieval-Augmented Code Generation (RACG) [67, 64, 68, 69] represents an emerging and promising area of research that harnesses the strengths of retrieval-based methods in conjunction with code generation techniques. Hayati et al. [64] introduced the RECODE model, which is the first to incorporate information retrieval techniques into the code generation task. Parvez et al. [68] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al. [67] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al. [69] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al. [70] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. the APIRetriever module of CodeGenAPI and our approach both focus on private API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
Code Completion Evaluation. LLM-based code completion and program synthesis methods provide unprecedented convenience, however researcher also question the actual performance under the mirage from different perspectives [50, 51, 34, 52, 16]. Yuan et al. evaluated 10 open-source instructed LLMs on four representative code comprehension and generation tasks [50]. Liu et al. proposed EvalPlus, a code synthesis benchmarking framework, to rigorously evaluate the functional correctness of LLM-synthesized code in function level [51] . In a more challenging code generation scenario, i.e., class-level code generation, Du et al. manually constructed the first class-level code generation benchmark ClassEval of 100 class-level Python code generation tasks [34]. In addition, Ouyang et al. conducted an empirical study to demonstrate that non-determinism in code generation tasks [52]. Yu et al. proposed a benchmark named CoderEval, which can be used to assess the performance of models against pragmatic code generation beyond just generating standalone functions [16].
8. Conclusion
Code generation within code repositories often necessitates comprehensive knowledge of the entire repository context. Existing techniques often lack such a holistic understanding of code repository knowledge, which is crucial for effective code generation. In this paper, we introduce the AAA Code Generation framework, which is designed to construct, mine, and unify three essential types of code repository knowledge: local-aware, global-aware, and third-party-library aware knowledge. To validate the effectiveness of these knowledge types, we conducted benchmark experiments and identified optimal configurations for their utilization. Our experiments demonstrate that our framework significantly enhances the ability of Language Model Models (LLMs) to fully utilize local, global, and third-party library information. This results in the generation of concise, high-quality code. In the future, we plan to expand the applicability of our framework to multiple programming languages. Furthermore, we aim to enhance our knowledge utilization methods to enable LLMs to become more aware of various types of code knowledge within code repositories.